In going from one city to another,a car travels 75 km north west and 20 km east. The approximate distance b/w two cities is?
Answer by sachi(548)

In going from one city to another,a car travels 75 km north west and 20 km east.
then it is clear that the car returns towards ist city
The approximate distance b/w two cities =75-20=55 km